# Crash Landing onto “you”: Untethered Soft Aerial Robots for Safe Environmental Interaction, Sensing, and Perching

### v1

@@
There are various desired capabilities to create aerial forest-traversing robots capable of monitoring both biological and abiotic data. The features range from multi-functionality, robustness, and adaptability. These robots have to weather turbulent winds and various obstacles such as forest flora and wildlife thus amplifying the complexity of operating in such uncertain environments. The key for successful data collection is the flexibility to intermittently move from tree-to-tree, in order to perch at vantage locations for elongated time. This effort to perch not only reduces the disturbance caused by multi-rotor systems during data collection, but also allows the system to rest and recharge for longer outdoor missions. Current systems feature the addition of perching modules that increase the aerial robots’ weight and reduce the drone’s overall endurance. Thus in our work, the key questions currently studied are: “How do we develop a single robot capable of metamorphosing its body for multi-modal flight and dynamic perching?”“How do we detect and land on perchable objects robustly and dynamically?”, and “What important spatial-temporal data is important for us to collect?”
@@
We have seen numerous bio-inspired principles as the bedrock of innovative solutions for aerial perching robots [4, 5]. Similar to grasping and gripping principles, it has ranged from fibre-based dry adhesives like gecko-inspired solutions, to deployable microspines, tensile perching like spiders, and avian-inspired grippers [6]. For forest environments, dry adhesives often struggle because of the roughness of the tree branches and deployable microspines are often paired with another grasping mechanism to enhance gripping quality. Tensile perching and avian-inspired grippers, on-the-other hand, require additional modules plugged onto a commercial unmanned aerial vehicle, thus increasing the overall weight of the robot, reducing the possible payload it could carry, and therefore compromising mission endurance.
@@
With the research objective centralized around the development of a metamorphic robot body and its interaction with different environments, one exciting strategy to develop aerial robots is through materials-based approaches seen in soft robotics [7]. Soft robotic methodologies open pathways to develop aerial perching robots that are imbued with soft robotic components, such as soft structures, sensing, and actuation [8, 7]. The goal would not be to create a completely soft-bodied aerial robot, because classically soft systems still suffer from slow actuation speeds, but use a combination of flexible, rigid and soft materials with encoded mechanical shape-changing and sensing capabilities instead. I believe this design methodology will lead to aerial robot designs that can manipulate and sense its shape for shared functionality, but are still safe and robust enough to interact and transition between flight and perching.
@@ II Future Work
Recently, different versions of morphing arm aerial robots (MAARs) have been developed, showcasing their ability in enabling adjustable flight dynamics and facilitating slow descent vertical perching [17, 18, 16]. Although these works establish an encouraging benchmark, I aim to extend the capabilities of these systems towards successful field missions by:

### v2

@@
There are various desired capabilities to create aerial forest-traversing robots capable of monitoring both biological and abiotic data. The features range from multi-functionality, robustness, and adaptability. These robots have to weather turbulent winds and various obstacles such as forest flora and wildlife thus amplifying the complexity of operating in such uncertain environments. The key for successful data collection is the flexibility to intermittently move from tree-to-tree, in order to perch at vantage locations for elongated time. This effort to perch not only reduces the disturbance caused by multi-rotor systems during data collection, but also allows the system to rest and recharge for longer outdoor missions. Current systems feature the addition of perching modules that increase the robots’ weight and reduce its overall endurance. Thus in our work, the key questions studied are: “How do we develop a single robot capable of metamorphosing its body for multi-modal flight and dynamic perching?”“How do we detect and land on perchable objects robustly and dynamically?”, and “What important spatial-temporal data is important for us to collect?”
@@
We have seen numerous bio-inspired principles as the bedrock of innovative solutions for aerial perching robots [4, 5, 6, 7, 8]. Similar to grasping and gripping principles, it has ranged from fibre-based dry adhesives like gecko-inspired solutions, to deployable microspines, tensile perching like spiders, and avian-inspired grippers [9]. For forest environments, dry adhesives often struggle because of the roughness of the tree branches and deployable microspines are often paired with another grasping mechanism to enhance gripping quality. Tensile perching and avian-inspired grippers, on-the-other hand, require additional modules plugged onto a commercial unmanned aerial vehicle, thus increasing the overall weight of the robot, reducing the possible payload it could carry, and therefore compromising mission endurance.
@@
With the research objective centralized around the development of a metamorphic robot body and its interaction with different environments, one exciting strategy to develop aerial robots is through materials-based approaches seen in soft robotics [10, 11]. Soft robotic methodologies open pathways to develop aerial perching robots that are imbued with soft robotic components, such as soft structures, sensing, and actuation [12, 11]. The goal would not be to create a completely soft-bodied aerial robot, because classically soft systems still suffer from slow actuation speeds, but use a combination of flexible, rigid and soft materials with encoded mechanical shape-changing and sensing capabilities instead. I believe this design methodology will lead to aerial robot designs that can manipulate and sense its shape for shared functionality, but are still safe and robust enough to interact and transition between flight and perching.
@@ II Future Work
Recently, different versions of morphing arm aerial robots (MAARs) have been developed, showcasing their ability in enabling adjustable flight dynamics and facilitating horizontal or vertical perching [21, 22, 20]. Although these works establish an encouraging benchmark, I aim to extend the capabilities of these systems towards successful field missions by: